RFBG Commissioning with Adaptive Secondary - 28 March 2010 UT
Observer: JHeidt, FBian (LBTO)
Telescope Operator: DHuerta
Instrument Support:: TShih
Telescope Support: JUrban
Adaptive Support: GBrusa (LBTO)
Weather was photometric, but only with moderate to poor seeing conditions
Tried to get rid of collimation problem with GS off x-axis by a change of
WFS_pupil_direction from -1 to 1
. Didn't help.
=> Doug sent Issue-Track 2637 about the X-pupil shift
Did a set of 50+ stars for a new pointing model using John's gopoint script
Did a couple of Persson stds in J and H to measure IRTC's zeropoint during the last hours.
WFS had problems to converge (sometimes diverged, sometimes limits hit).
Single frame names refer to: irtc.20100328.xxxxxx.fits
Cube names refer to: irtc.20100328_00xxx_cube.fit
Plan for the night: Do a large set of transformation data, collect field aberration measurements (40" tip)
If seeing is not so good: do pointing model observations, if seeing is bad: IRTC ZP-measurements in J and H (its photometric)
Check also if GCS-fix for X-axis collimation problem is OK
Cubes 1-20 do not have "irtc" as prefix, sorry.
01:45 Open, clear (photometric), T_out = -5C, T_Mirror = -3.1C, Hum = 45%, a tiny bit of wind
Initial Pointing and Collimation Adjustment
Each night in this RFBG commissioning run, we go to a bright star (Pollux) at Elevation 80 degrees. Then we adjust the vertex tip-tilt (RX/RY) of the secondary to get rid of the coma
that we see in the extrafocal pupil on the IRTC.
01:50 pointto Pollux gs=0, TelM
= track, PAMode=no, no star, adjust to yesterday values (CA=-77, IE=-48), EL = 78, OK, star there, adjust *CA=-71 ,IE=-49*
DX M2 Global X = 2.5, Y = -3.63, Z = 1.9, RX = -360, RY = -120
, looks OK, no obvious coma, possibly
a bit of astigmatism, no adjustment necessary
01:56 pointto M5_1210 gs=0, TelM
= track , EL = 26, AZ = 170, at very-low elevation to adjust AZ-zeropoint, pointing OK, star found, Adjust IA = -383
, keep IE=-49, CA = -71. Coma, Adjust global offsets to:
DX M2 Global X = 2.2, Y = -2.7, Z = 2.5, RX = -490, RY = -135
02:18 pointto BS9134, GS=0, TelM
=active, check for collimation/seeing, takes a little while, looks OK, guider reports ~1.3",
RMS@WFE ~150, seeing unfortunately varying a bit between 1.2 and 1.8"
Check if GCS-fix for X-axis collimation problem is OK
02:34 pointto BS9134, GS=9, TelM
=active, PA=INDEF, (sets PA=302.34), RMS@WFS 150...300, star looks round
02:43 pointto BS9134, GS=9, TelM
=active, PA=240, check for x-axis collimation, doesn't work
, GS moves badly out of focus
Doug tries to find out what's wrong
Ask David to clear active optics, and focussing by eye...
02:34 pointto BS9134, GS=0, TelM
=active, to do collimation, OK, works
Pointing model observations (part1)
T_out = -5.0, T_mirror = -4.1 (i.e. almost in therm. equilibrium)
Take pair of TRACK mode images rotated 180degs apart
03:17 pointto BS9134, GS=9, TelM
=track, PA=INDEF, (sets PA=302.34), take image 031917 (10sec)
03:21 pointto BS9134, GS=9, TelM
=track, PA=122.34), take image 032231 (10sec)
Turned on pointing logging, Filename is 2010march28UT
03:38 pointto M5_0347, GS=0, TelM
=track, check pointing, OK, star there, CA = -71, IE = -48
Lets see if and how it works:
03:41 started gopoint (gridID=50, agwNum=2, exptime=0.1, Nimages=100, gologfile=gopointmar282010)
stopped since script doesn't allow for agwNum=2
03:49 pointto BS9134, GS=0, TelM
=active, let collimate, RMS@WFS ~200, OK, guider reports 1.6-2"
03:59 started gopoint again (gridID=50, agwNum=indef, exptime=0.1, Nimages=100, gologfile=gopointmar282010)
Modified script to allow for AGW2
04:08 started gopoint again (gridID=50, agwNum=indef, exptime=0.1, Nimages=100, gologfile=gopointmar282010),
seems to run now.
05:15 Need to reset SM after 23 stars (by Guido)
05:18 pointto BS9134, GS=0, TelM
=active, let collimate
05:25 Need to reset SM again (by Guido, who has to call Italy to get rid of the problem)
In the meantime Doug found a possible solution to get rid of the X-axis collimation problem. Tony Edgin was contacted to
make a parameter change for the active optics to work properly.
06:10 Tony Edgin incorporated the parameter change requested by Doug, will be tested as soon as the shell is working again.
07:05 Dome closed, still problems with shell, Guido needs to make a few tests.
07:40 Opening again, David cleared active optics
Check if parameter change solves the X-axis collimation problem
08:00 pointto BS9153, GS=0, TelM
=active, lets collimate, star not found
08:04 pointto M5_1396, GS=0, TelM
=track, check pointing, OK, star found, CA=-55, IE=-40
08:08 pointto BS9153, GS=0, TelM
=active, lets collimate, seeing has improved, guider reports ~1", RMS@WFS below 100
08:14 pointto BS9153, GS=10, TelM
=active, PA=INDEF (sets PA=187.83), RMS@WFS below 500, 1"
wind a little high, David has to close the vent doors
08:22 pointto BS9153, GS=10, TelM
=active, PA=127.83, check for X-axis collimation, no improvement
Restart GCS (to make shure that parameter changes are active)
08:34 pointto BS9153, GS=10, TelM
=acquire, PA=INDEF, for rough focus adjustment
08:15 pointto BS9153, GS=10, TelM
=active, PA=INDEF (sets PA=187.83), collimation, Ok, RMS@WFS ~400
08:41 pointto BS9153, GS=10, TelM
=active, PA=127.83, check for X-axis collimation, does not
Pointing model observations (part2)
T_out = -5.6, T_mirror = -5.5 (i.e. in perfect therm. equilibrium)
08:48 pointto BS9153, GS=0, TelM
=active, lets collimate, guider reports ~ 1.3", RMS@WFS below 200
Pointing logging is still on, Filename is still 2010march28UT
08:53 started gopoint again (gridID=74, agwNum=indef, exptime=0.1, Nimages=100, gologfile=gopointmar282010),
seems to run now. Stopped at Grid-Number 88 (I started with 50 and it runs only to that point..
09:54 started gopoint again (gridID=1, agwNum=indef, exptime=0.1, Nimages=100, gologfile=gopointmar282010), stopped after grid position 20 stars. In total ~55 stars have been observed.
10:51 pointto BS9177, GS=0, TelM
=acquire, to focus by eye via David.
11:06 pointto BS9177, GS=0, TelM
=active, to collimate, not easy, seeing varies strongly between 1.5 and 3" => no chance for field aberration measurements => do some ZP-measurements in J and H (its photometric)
IRTC J and H ZP's
Always 2sec single and 10*2sec cube
11:16 pointto BS9177, GS=19, TelM
=active, EL=60, M1 reached limits, stopped guiding, cleared active optics,
redo pointing, active optics again diverges
11:38 pointto BS9164, GS=0, TelM
=active, OK, converges
11:38 pointto BS9164, GS=8, TelM
H, 114908 cube 3
J, 115014 cube 4
H, 115441 cube 5, offset by 15" for sky subtraction
J, 115532 cube 6, offset by 15" for sky subtraction
11:56 pointto BS9162, GS=11, TelM
H, 120303 cube 7
J, 120350 cube 8
Something is wrong, hit again limits, again cleared optics
12:06 pointto BS9160, GS=19, TelM
=active, hit once more limits again????????
H, 121916, cube 9
J, 122004, cube 10
12:25 pointto BS9158, GS=4, TelM
H, 122844, cube 11
J, 122929, cube 12
12:30 pointto BS9156, GS=9, TelM
H, 123332, cube 13
J, 123415, cube 14
12:35 pointto BS9155, GS=8, TelM
H, 123818, cube 15
J, 123903, cube 16
12:40 pointto BS9153, GS=10, TelM
H, 124202, cube 17
J, 124247, cube 18
12:43 pointto BS9152, GS=1, TelM
H, 124612, cube 19, didn't find guid star, stop guiding, blind image
J, erroneously not done
H, 124933, cube 21, offset by 15" for sky subtraction
J, 125019, cube 22, offset by 15" for sky subtraction
Turned off pointing logging a bit late, sorry...
12:52 Too bright now, T_out = -6.0C, T_Mirror = -5.9, Hum = 40%, photometric
- 28 Mar 2010